Question 6 (1 point Fill in the following contingency table and find the number of students who both do not watch comedies AND do not watch dramas. watch dramas do not watch dramas Total 36 Students 75 56 watch comedies do not watch comedies Total 64 Select the correct answer below: 067 ● 131 25 31 061
Question 7 (1 point Suppose you draw two cards from a 52-card deck without replacement. What is the probability that you draw 2 aces? Round to the nearest thousandths. Provide your answer below:
Question 8 t1 poln A bag contains 7 RED beads, 11 BLUE beads, and 6 GREEN beads. If a single bead is picked random, what is the probability that the bead is BLUE or GREEN? Select the correct answer below: 24 18 24 7 24 24 6 24 13 24
Question 9 (1 point) Using the following set of data, find the sample standard deviation: 6, 10, 9, 13, 5,5 Select the correct answer below: 5.2 3.2 9.2 6.2 1.2 Previous Page Next Page Page
Answer #1

6.

Student watch dramas do not watch dramas Total
watch comedies 36 39 75
do not watch comedies 31 25 56
Total 67 64 131

Number of students who both do not watch comedies and dramas is 25

7.

There are 4 aces in a deck of 52 cards.

Number of ways to draw 2 cards from 52 cards = 52C2 = 1326

Number of ways to draw 2 aces from 4 = 4C2 = 6

Probability to draw both aces = 6 / 1326 = 0.004524887

8.

Total number of blue or green beads = 11 + 6 = 17

Total number of beads = 7 + 11 + 6 = 24

Probability that the selected bead is blue or green = 17/24

9.

Mean = (6 + 10+9+13+5+5) / 6 = 8

Sample standard deviation = sqrt[ ((6 - 8)2 + (10 - 8)2 + (9 - 8)2 + (13 - 8)2 + (5 - 8)2 + (5 - 8)2 ) / (6 - 1)]

= 3.2